Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
mwas - 21:16 Wednesday 22 May 2024 (64328) Print this report
Range decrease to 45Mpc because of etalon

Figure 1. After the day mostly unlocked yesterday there was a large transient on etalon temperature control. This resulted in a decrease in optical gain, decrease in BNS range, and the DIFFp TX error signal based on the DARM line moving away from zero. It underlines the importance of implementing the scheme tested at the end of O3 to keep the etalon control stable during unlocks.

Images attached to this report
Comments to this report:
mantovani - 15:36 Thursday 23 May 2024 (64332) Print this report

I think it is hard to address all the lost MPc to etalon. I fear that an other drift (not clear of which parameters) is adding on the top.

As it is visible for NI te < 20.08, the incriminated drift, the behavior of BNS and DIFFp tx has changed behavior and it can point to an other monotonic drift adding on the top of the Etalon drift.

In any case improving the stability of Etalon will better disantangle the other drifts of the machine helping the understanding

Images attached to this comment
Search Help